Q.Choose the correct answer.
Q. The equation u_xx 〖-4u〗_xy=0 is
)parabolic
b)hyperbolic
c)Elliptic
d) None of above
1
Expert's answer
2019-07-10T12:09:32-0400

The equation is Au_xx+2Bu_xy+Cu_yy. In our case A=1, 2B=-4, B=-2, C=0 (there is no u_yy in the equation). D=B^2-AC=(-2)^2-1*0=4>0, which indicates that the type of the equation is hyperbolic. The answer is b) hyperbolic.